# Search-based DNN Testing and Retraining with GAN-enhanced Simulations

### v1

@@ Abstract
In safety-critical systems (e.g., autonomous vehicles and robots), Deep Neural Networks (DNNs) are becoming a key component for computer vision tasks, particularly semantic segmentation. Further, since the DNN behavior cannot be assessed through code inspection and analysis, test automation has become an essential activity to gain confidence in the reliability of DNNs. Unfortunately, state-of-the-art automated testing solutions largely rely on simulators, whose fidelity is always imperfect, thus affecting the validity of test results. To address such limitations, we propose to combine meta-heuristic search, used to explore the input space using simulators, with Generative Adversarial Networks (GANs), to transform the data generated by simulators into realistic input images. Such images can be used both to assess the DNN performance and to retrain the DNN more effectively. We applied our approach to a state-of-the-art DNN performing semantic segmentation and demonstrated that it outperforms a state-of-the-art GAN-based testing solution and several baselines. Specifically, it leads to the largest number of diverse images leading to the worst DNN performance. Further, the images generated with our approach, lead to the highest improvement in DNN performance when used for retraining. In conclusion, we suggest to always integrate GAN components when performing search-driven, simulator-based testing.
@@
Deep Neural Networks (DNNs) are an effective solution to automate tasks that require vision capabilities. For this reason, several companies developing cyber-physical systems (CPS), are making large R&D investments to rely on DNNs to automate tasks that are currently manual. Examples in the automotive sector include driving assistance systems (e.g., obstacles detection, traffic sign detection, autonomous driving) and interior car monitoring (e.g., drowsiness detection or unsupervised child detection). In other sectors, key developments are observed in robotics and space, for example to develop autonomous systems capable of exploring caves and planets’ surfaces.
@@
When a DNN that processes camera images is used to drive a CPS, we must gain sufficient confidence that it can adequately respond to all foreseeable inputs. To achieve this objective, research has focused on combining meta-heuristic search with simulators [25, 26, 21, 19, 42, 62]. The rationale is that, under the assumption that simulators can accurately create images corresponding to a realistic scene, meta-heuristic search enables to cost-effectively search for inputs leading to failures. What enables test automation is the fact that simulators, in addition to generate images according to the parameters selected by the search algorithm, can be used to automatically assess the quality of the DNN output for each generated input. For example, the performance of steering angle DNNs can be assessed by relying on the distance between the ego vehicle and the lane separator line, both provided by the simulator; indeed, the closer the vehicle gets to the separator line, the more likely it will cross the lane. Although such search-based approaches effectively generate images leading to DNN failures, the validity of the assessed DNN reliability highly depends on the simulator fidelity. For example, recent work has demonstrated that the fidelity gap of simulators leads to different reliability results in simulation-based and real-world testing [26, 46].
@@
The limited fidelity of a simulator might thus hinder the detection of DNN shortcomings. For example, the road landscapes generated by the AirSim simulator [44], which resemble residential areas in the USA, are very different from the landscape of European cities and towns, which can be the working environment of road scene segmentation DNNs (e.g., detecting the position of other cars) that can be successfully trained using the CityScape dataset [13]. Based on our interactions with industry partners in the space and automotive sector [17, 18, 4], we noticed that high-fidelity simulators are lacking for many environments in which DNNs can be applied, including for example car interiors (e.g., for in-cabin monitoring DNNs) and space landscapes (e.g., to drive Mars and Lunar rovers).
@@
To accurately test DNNs, ensuring that we explore the input space relying on high-fidelity input images, we propose DNN testing with GAN-enhanced simulations and search (DESIGNATE). We rely on meta-heuristic search to effectively drive a simulator towards the generation of diverse inputs leading to failures but, instead of testing the DNN with low-fidelity simulator images, we leverage Generative Adversarial Networks (GANs) to transform the simulator output into a realistic image that closely resembles real-world data distributions. GANs are a promising candidate to address the simulator fidelity gap because they have shown to be effective in image-to-image translation, that is mapping images from one domain (e.g., sketches) to images in another (e.g., photographs) [29, 51]. In testing, GANs have been mainly adopted to introduce realistic perturbations in input images [58, 32, 52] and to estimate sensory data [56, 46], but DESIGNATE is the first technique integrating GANs into search-based testing with simulators.
@@
We assume that, because the realistic images are generated from the simulated ones, both the simulator and realistic images should share the same ground truth (i.e., the DNN should generate similar outputs from them), thus enabling test automation by assessing the DNN output based on the information provided by the simulator. Further, realistic images leading to DNN failures can be used to retrain a DNN to improve its reliability in the field, something difficult to achieve with simulator images due to their possible low fidelity. For example, DESIGNATE can test a road scene segmentation DNN by relying on AirSim to generate landscapes of city roads with traffic and buildings, and then leverage a GAN to transform those landscapes into realistic images with European cities.
@@
Inspired by work on DNN testing [26, 42], we rely on genetic algorithms to cost-effectively identify failure-inducing images with GANs; our objective is to identify simulator parameters leading to simulator outputs (e.g., an image of a simulated city road) that, after being transformed into realistic images using a GAN, lead to DNN failures. Since we ideally aim at identifying all the situations in which the DNN fails, to both identify failures and maximize input diversity, we combine two fitness functions: The first measures the accuracy of the DNN output. The second measures the similarity across generated realistic images. To test a DNN with diverse scenes being depicted in the input images, we are interested in generating images that differ in terms of the items being displayed and not just a few, scattered, unmatched pixels. For this reason, we introduce a feature-based distance metric that leverages transfer learning [4].
@@
We conducted an empirical investigation with a state-of-the-art, third-party DNN for road segmentation. Our results show that DESIGNATE outperforms TACTIC [32], a state-of-the-art DNN testing approach, and several baselines including random search, single objective search, and multi-objective search without GANs. DESIGNATE generates inputs leading to the worst DNN reliability and it maximizes the number of distinct failure scenarios. Further, we demonstrate that the inputs generated by DESIGNATE better support DNN retraining, compared to competing approaches, with an increase in DNN accuracy by 8 percentage points. Based on our results, we conclude that GANs should be combined with simulators when testing and improving vision-based DNNs.
@@ II-C Simulated Environments
Simulators replicate real-world scenarios, allowing DNNs to be tested in a controlled and reproducible environment. In our work, we rely on AirSim, an open-source simulator for autonomous vehicles and drones developed by Microsoft [44]. It provides a realistic environment for testing and training AI algorithms using a variety of sensors, including cameras, lidars, and GPS.
AirSim can be controlled through a dedicated API, which enables its use for DNN testing. In our context, we rely on the features for controlling the positioning of the ego vehicle in the simulated environment and generating a picture of the landscape in front of the ego vehicle. Further, for each generated image, AirSim provides a semantic segmentation image where each pixel’s color directly corresponds to a specific category or class, including roads, buildings, vehicles, pedestrians, or sky. We use these images as ground truth to assess the reliability of DNNs in road scene segmentation.
@@ III The DESIGNATE Approach
GAN-based input generation component. It receives as input the configuration parameter values to be used to generate, with a simulator, scene data (i.e., an image and the corresponding segmentation map). The scene data is then used to generate a realistic image with a GAN trained on a specific reference domain. In our empirical assessment, we use the Pix2PixHD GAN trained on the CityScape dataset to generate realistic road scenes from the segmentation map produced by the simulator.
@@ III-B1 Accuracy fitness
Determining out-of-domain inputs depends on the specific case study. In our empirical study, for example, we observe out-of-domain inputs when the ego vehicle is positioned next to another vehicle and perpendicular to it (e.g., facing the other vehicle’s door), which is not interesting for our testing purposes since it is very easy to make a segmentation DNN fail in such case. Another case is that of images with no cars, which should be excluded because we are interested in testing the DNN with images exhibiting traffic. Consequently, we consider an image to be relevant if the proportion of pixels belonging to cars (i.e., ) in the input image is within a given range:
We set the upper bound to to ensure that the car does not cover most of the image. We set the lower bound above zero to ensure that the image contains at least one car. The variable carPixelsProportion is computed on the segmentation ground truth.
@@
After sorting the generated image pairs based on their , it is possible, through visual inspection, to determine a threshold () above which the realistic images do not correctly capture the ground truth (e.g., a car is missing where expected by the ground truth). During testing, we thus consider a realistic image to be relevant if . For the segmentation DNN considered in our empirical assessment:
and we selected as the third percentile of the distribution.
In the segmentation DNN considered in our empirical assessment, is therefore computed as follows:
@@ III-C Search algorithm
We aim to address a multi-objective problem, generating realistic images leading to minimal DNN accuracy and having minimal similarity among each other. We therefore rely on the NSGA-II algorithm, which demonstrated its effectiveness in multiple multi-objective software testing problems. However, inspired by related work aiming at maximizing the diversity among the generated individuals [42], we extend NSGA-II with an archive to store the best (non-dominated) individuals encountered during the search process. The archive plays a critical role in preventing the genetic algorithm from cycling, a phenomenon where the population moves from one area of the solution space to another and then back again without exploring other areas. By retaining the best individuals in the archive, and assigning high fitness values to individuals that are similar to the already generated ones, the algorithm ensures that these areas are not revisited, facilitating a more efficient exploration of the solution space.
@@
Algorithm 2 presents the pseudocode of the NSGA-II algorithm (with archive) employed in DESIGNATE. After creating an empty archive (Line 3), since the initial selection of seeds may affect the search results, inspired by our previous work [3], we generate multiple ( in our experiments) populations of randomly selected individuals (Line 4) and we select the population as the one with the individual having the lowest (Line 6). All the individuals in are then used to update the archive based on Algorithm 1 (Line 7). In practice, the archive is populated with individuals whose distance is above and lead to worst accuracy.
@@ IV Empirical Evaluation
RQ1: How does DESIGNATE compare to alternative DNN testing solutions in terms of test effectiveness? Alternatives include approaches generating test images by relying on GANs (but not simulators), random baselines, and alternative implementations of DESIGNATE (ablation study). We discuss what approach leads to inputs leading to the lowest DNN accuracy and having the highest diversity.

### v2

@@ Abstract
In safety-critical systems (e.g., autonomous vehicles and robots), Deep Neural Networks (DNNs) are becoming a key component for computer vision tasks, particularly semantic segmentation. Further, since DNN behavior cannot be assessed through code inspection and analysis, test automation has become an essential activity to gain confidence in the reliability of DNNs. Unfortunately, state-of-the-art automated testing solutions largely rely on simulators, whose fidelity is always imperfect, thus affecting the validity of test results. To address such limitations, we propose to combine meta-heuristic search, used to explore the input space using simulators, with Generative Adversarial Networks (GANs), to transform the data generated by simulators into realistic input images. Such images can be used both to assess the DNN accuracy and to retrain the DNN more effectively. We applied our approach to a state-of-the-art DNN performing semantic segmentation, in two different case studies, and demonstrated that it outperforms a state-of-the-art GAN-based testing solution and several other baselines. Specifically, it leads to the largest number of diverse images leading to the worst DNN accuracy. Further, the images generated with our approach, lead to the highest improvement in DNN accuracy when used for retraining. In conclusion, we suggest to always integrate a trained GAN to transform test inputs when performing search-driven, simulator-based testing.
@@
Deep Neural Networks (DNNs) are an effective solution to automate tasks that require vision capabilities. For this reason, several companies developing cyber-physical systems (CPS), are making large R&D investments to rely on DNNs to automate tasks that are currently manual. Examples in the automotive sector include driving assistance systems (e.g., obstacles detection, traffic sign detection, autonomous driving) and interior car monitoring (e.g., drowsiness detection or unsupervised child detection). In other sectors, key developments are observed in robotics and space, for example to develop autonomous vehicles capable of exploring caves and planets’ surfaces.
@@
When a DNN that processes camera images is used to drive a CPS, we must gain sufficient confidence that it can adequately respond to all foreseeable inputs. To achieve this objective, research has focused on combining meta-heuristic search with simulators [26, 27, 22, 20, 43, 68]. The rationale is that, under the assumption that simulators can accurately create images corresponding to a realistic scene, meta-heuristic search enables the cost-effective search of inputs leading to failures. What enables test automation is the fact that simulators, in addition to generate images according to the parameters selected by the search algorithm, can be used to automatically assess the quality of the DNN output for each generated input. For example, the performance of steering angle DNNs can be assessed by relying on the distance between the ego vehicle and the lane separator line, both provided by the simulator; indeed, the closer the vehicle gets to the separator line, the more likely it will cross the lane. Although such search-based approaches effectively generate images leading to DNN failures, the validity of such DNN assessment highly depends on the simulator fidelity. For example, recent work has demonstrated that the fidelity gap of simulators leads to different reliability results in simulation-based and real-world testing [27, 49].
@@
The limited fidelity of a simulator may thus hinder the detection of DNN shortcomings. For example, the road landscapes generated by the AirSim simulator [46], which resemble residential areas in the USA, are very different from the landscape of European cities and towns. Based on our interactions with industry partners in the space and automotive sectors [18, 19, 5], we noticed that high-fidelity simulators are lacking for many environments in which DNNs can be applied, including for example car interiors (e.g., for in-cabin monitoring DNNs) and space landscapes (e.g., to drive Mars and Lunar rovers).
@@
To accurately test DNNs, ensuring that we explore the input space relying on realistic input images, we propose DNN testing with GAN-enhanced simulations and search (DESIGNATE). We rely on meta-heuristic search to effectively drive a simulator towards the generation of diverse inputs leading to failures but, instead of testing the DNN with simulator images, we leverage Generative Adversarial Networks (GANs) to transform the simulator output into a more realistic image that resembles real-world data distributions. GANs are a promising candidate to address the simulator fidelity gap because they have shown to be effective in image-to-image translation, that is mapping images from one domain (e.g., sketches) to images in another (e.g., photographs) [30, 56]. For testing, GANs have been mainly adopted to introduce realistic perturbations in input images [64, 33, 57] and to estimate sensory data [62, 49]. DESIGNATE is the first technique integrating GANs into search-based testing with simulators.
@@
We assume that, because the realistic images are generated from the simulated ones, both the simulator and realistic images share the same ground truth (i.e., the DNN should generate similar outputs from them), thus enabling test automation by assessing the DNN output based on the information provided by the simulator. Further, realistic images leading to DNN failures can be used to retrain a DNN to improve its reliability in the field, something which may be more difficult to achieve with simulator images because of on their fidelity. However, to address this challenge, DESIGNATE can test a road scene segmentation DNN by relying on AirSim to generate landscapes of city roads with traffic and buildings, and then leverage a GAN to transform those landscapes into realistic images with European cities.
@@
Inspired by work on DNN testing [27, 43], we rely on genetic algorithms to cost-effectively identify failure-inducing images with GANs; our objective is to identify simulator parameters leading to simulator outputs (e.g., an image of a simulated city road) that, after being transformed into realistic images using a GAN, lead to DNN failures. Since we ideally aim at identifying all the situations in which the DNN fails, to both identify failures and maximize input diversity, we combine two fitness functions. The first one measures the accuracy of the DNN output whereas the second one measures the similarity across generated realistic images. To test a DNN with diverse scenes being depicted in the input images, we are interested in generating images that differ in terms of the items being displayed and not just a few, scattered, unmatched pixels. For this reason, we introduce a feature-based distance metric that leverages transfer learning [5].
@@
We conducted an empirical investigation with a state-of-the-art, third-party DNN for road segmentation. Our results show that DESIGNATE outperforms TACTIC [33] and DeepJanus [43], two state-of-the-art DNN testing approaches, and several baselines including random search, single objective search, and multi-objective search without GANs. DESIGNATE generates inputs leading to the worst DNN reliability and maximizes input diversity. Further, we demonstrate that the inputs generated by DESIGNATE better support DNN retraining, compared to competing approaches, with an increase in DNN accuracy of eight percentage points. ††margin: R3.6 In addition to our road segmentation case study, in the context of the Mars rover self-driving system, we also demonstrate the applicability and benefits of our approach to object detection on Mars’ surface, using the AI4MARS dataset [50] and a custom-built simulator leveraging Unreal Engine [16] and Airsim. Based on our results across both case studies, we conclude that GANs should be combined with simulators when testing and improving vision-based DNNs as it leads to more effective and diverse testing, as well as better retraining.
@@ II-C Simulated Environments
Simulators replicate real-world scenarios, allowing DNNs to be tested in a controlled and reproducible environment. In our work, we rely on two simulators:
II-C1 AirSim
AirSim is an open-source simulator for autonomous vehicles and drones developed by Microsoft [46]. It provides a realistic environment for testing and training AI algorithms using a variety of sensors, including cameras, lidars, and GPS. AirSim can be controlled through a dedicated API, which enables its use for DNN testing. In our context, we rely on the features for controlling the positioning of the ego vehicle in the simulated environment and generating a picture of the landscape in front of the ego vehicle. Further, for each generated image, AirSim provides a semantic segmentation image where each pixel’s color directly corresponds to a specific category or class, including roads, buildings, vehicles, pedestrians, or sky. We use these images as ground truth to assess the reliability of DNNs in road scene segmentation.
II-C2 MarsSim
††margin: R3.6Focusing on extraterrestrial object detection, specifically on the Martian surface, we built a simulator that generate images resembling the ones in the AI4MARS dataset by combining Unreal Engine with the the same plugins that Airsim uses. This simulator creates realistic images of Mars terrain, which, when enhanced with a GAN, enables the generation of diverse training data for testing DNNs in Mars exploration contexts. Including the Airsim API enables us to control the simulator by changing the position of the camera. To create an authentic Martian environment, we modeled objects that match those found in the AI4MARS dataset: rocks, bedrock, sand, and Martian soil, each designed with textures and colors reflective of actual Mars imagery. We also incorporated variations in object size, shape, and placement to reflect the irregular distribution of terrain features typically observed on Mars. Additionally, we introduced terrain undulations and slopes to simulate the topographical challenges of Mars, such as craters, cliffs, and valleys, which pose unique navigational challenges for autonomous rovers.
@@ III The DESIGNATE Approach
GAN-based input generation component. It receives as input the configuration parameter values to be used to generate, with a simulator, scene data (i.e., an image and the corresponding segmentation map). The scene data is then used to generate a realistic image with a GAN trained on a specific reference domain. For our automotive case study, we used the Pix2PixHD GAN trained on the CityScape dataset to generate realistic road scenes from the segmentation map produced by the simulator. ††margin: R3.6 For our Mars case study, we utilized the AI4MARS dataset to train the GAN, and the simulator was built using Unreal Engine and Airsim to generate Martian surface images.
@@ III-B1 Accuracy fitness
For the Martian case study, we consider all the classes when computing the accuracy. Therefore, we use the mean IoU () as . The for one image is obtained by averaging the IoU scores across all classes and is computed as follows:
@@
How we determine out-of-domain inputs depends on the specific case study. When relying on AirSim, for example, we observe out-of-domain inputs when the ego vehicle is positioned next to another vehicle and perpendicular to it (e.g., facing the other vehicle’s door), which is not interesting for our testing purposes since it is very easy to make a segmentation DNN fail in such case. Another case is that of images with no cars, which should be excluded because we are only interested in testing the DNN with images exhibiting traffic. Consequently, we consider an image to be relevant if the proportion of pixels belonging to cars (i.e., ) in the input image is within a given range:
The variable carPixelsProportion is computed on the segmentation ground truth. We set the upper bound to to ensure that the car does not cover most of the image, which would result in unrealistic scenarios for evaluating DNN performance. Indeed, images where carPixelsProportion exceeds 0.4 often involve close-ups or unusual viewpoints, such as a vehicle occupying most of the camera’s field of view—lidar-based sensors typically prevent autonomous cars from being in such situations. We set the lower bound above zero to ensure that the image contains at least one car. ††margin: R3.6 Similarly, for the Martian environment, we need to ensure that the sky does not cover most of the image and, therefore, we consider an image relevant if the proportion of pixels belonging to the sky (i.e., ) is within the range:
One may consider that is a relatively large proportion of the image but it is a realistic situation based on our discussions with ESA representatives.
@@
After sorting the generated image pairs based on their , it is possible, through visual inspection, to determine a threshold () above which the realistic images do not correctly capture the ground truth (e.g., a car is missing where expected based on the ground truth). We selected as the third percentile of the distribution. During testing, we thus consider a realistic image to be relevant if . Regarding the segmentation DNN for urban environments considered in our empirical assessment:
In contrast, in the case of the martian environment, we did not observe GAN-generated images that are less realistic than those of the simulator and therefore we do not rely on . We believe this is mainly due to our Mars simulator having limited fidelity and is thus representative of cases where limited effort was dedicated to simulator development111Please note that state-of-the-art realistic simulators for space exist (e.g., PANGU [48]), but they do not generate the ground truth labels required to train segmentation DNNs..
In the segmentation DNN considered in our empirical assessment with the urban case study, is therefore computed as follows:
For the martian case study, based on the above discussion, is instead computed as follows:
@@ III-C Search algorithm
We aim to address a multi-objective problem [4], generating realistic images leading to minimal DNN accuracy and having minimal similarity among each other. We therefore rely on the NSGA-II algorithm, which demonstrated its effectiveness in multiple multi-objective software testing problems. However, inspired by related work aiming at maximizing the diversity among the generated individuals [43], we extend NSGA-II with an archive to store the best (non-dominated) individuals encountered during the search process. The archive plays a critical role in preventing the genetic algorithm from cycling, a phenomenon where the population moves from one area of the solution space to another and then back again without exploring other areas. By retaining the best individuals in the archive, and assigning high fitness values to individuals that are similar to the already generated ones, the algorithm ensures that these areas are not revisited, facilitating a more efficient exploration of the solution space.
@@
Algorithm 2 presents the pseudocode of the NSGA-II algorithm (with archive) employed in DESIGNATE. After creating an empty archive (Line 3), since the initial selection of seeds may affect the search results and inspired by our previous work [3], we generate multiple ( in our experiments) populations of randomly selected individuals (Line 4) and we select the population as the one with the individual having the lowest (Line 6). ††margin: R3.1 This choice is based on empirical observations from prior studies [19], which demonstrated that using multiple seeds helps to reduce bias in the optimization process and ensures a broader exploration of the solution space. For instance, using five seeds strikes a balance between computational efficiency and the reliability of search results in related search-based testing tasks. All the individuals in are then used to update the archive based on Algorithm 1 (Line 7). In practice, the archive is populated with individuals whose distance is above and lead to worst accuracy.
@@ IV Empirical Evaluation
RQ1: How does DESIGNATE compare to alternative DNN testing solutions in terms of test effectiveness? Alternatives include approaches generating test images by relying on GANs (but not simulators), random baselines, and alternative implementations of our approach (ablation study). We discuss what approach leads to inputs leading to the lowest DNN accuracy and having the highest diversity.
@@
IV-A Subjects of the Study
As subjects of our study, we consider two distinct subjects: 1) The DeeplabV3 DNN [10] from the Gluoncv library [23], which is a state-of-the-art semantic segmentation DNN that classifies each pixel image into a predefined class. This DNN focuses on urban environments relying on the Cityscapes dataset [45]. 2) A DeeplabV3 DNN trained for object detection on Mars, utilizing the AI4MARS dataset [50]. This DNN implements a key task for the navigation of Mars rovers; specifically, it detects objects such as rocks, sand, and bedrocks. This subject was developed in the context of a project with the European Space Agency.
For the urban environment case, DeeplabV3 was initially trained by its developers on the Cityscapes dataset, which contains a variety of urban street scenes from different European cities. The dataset consists of a variety of images captured at a resolution of 1024 by 2048 pixels in the streets of 50 different European cities, primarily in Germany. Each real-world image comes with a segmentation mask representing its ground truth. The DeeplabV3 training set consists of 2975 images; however, its test set comes without ground truth and, therefore, for testing, we rely on the Cityscapes validation set (500 images). ††margin: R3.3 The average metric (see Section III) is particularly high: 71% (min=00%, first-quartile=62%, third-quartile=91%, max=97%).
For the Mars navigation subject, we trained the DNN on the AI4MARS dataset, which provides annotated images of the Martian surface. The simulator for generating synthetic Martian images was built using Unreal Engine [16] with the same plugins used by AirSim. The simulator was configured to simulate realistic Martian terrain and lighting conditions. The Airsim plugins are added to: simulate Vehicle dynamics (including interaction with sand and soil), add LIDAR for object detection, provide Python API for real-time control of vehicle, enabling communication between our search algorithm and the simulated vehicle. For segmentation on the AI4MARS dataset, DeepLabV3 achieved an average mIoU of 49% (min=0%, first-quartile=1%, third-quartile=77%, max=99%).
IV-B Assessed techniques
In our experiments, we assess a DESIGNATE prototype, that relies on ††margin: R2.1 the original pix2pixHD GAN [38] and AirSim 1.8.1 [37] for the urban case study, and Unreal Engine 5 for the Martian case study. We also extended the NSGA-II algorithm provided by the Pymoo library [7].
To address our RQs, we compare DESIGNATE with a random baseline, two variants of DESIGNATE (for the purpose of performing ablation studies), ††margin: R2.1 TACTIC [33], and DeepJanus [43]. We selected DeepJanus because it may lead to better results if a DNN tends to fail where model behaviour changes. For DeepJanus, in addition to the original implementation provided by its authors, we developed an extended version that integrates a GAN, as described below.
Our random baseline relies only on the GAN-based input generation component; specifically, it generates realistic images from simulator segmentation maps generated using randomly selected values for the simulator parameters (e.g., position of the car in Airsim). For brevity, we refer to the images from our random baseline as random realistic images. We generate as many random images as DESIGNATE (i.e., ).